# Mitigating Biases in Surgical Operating Rooms with Geometry

### v1

@@ 1 Introduction
Figure 1: RGB image excerpts and overlayed saliency maps generated with GradCAM [13] on the simulated datasets 4D-OR [12] and $M$-OR [17]. 4D-OR’s limited realism and variety allow CNNs to identify individuals solely by their heads and shoes. In more realistic OR settings like $M$-OR, these features become less useful due to more homogenous attire. These discrepancies between different clinical environments can impede generalization.
@@ 2 Methods
We trained ResNet-50 models for person re-identification on two surgical datasets: the simulated 4D-OR [12] and the more realistic $M$-OR [17]. Using GradCAM [13], we generated class activation maps to visualize which spatial regions the networks relied upon for identity discrimination. Models are trained on a supervised classification task of the individuals’ known identity.
@@
We evaluate our approach on two surgical datasets: the simulated 4D-OR [12] containing 5 individuals across 10 procedures, and the more authentic $M$-OR [17] containing 13 individuals across 11 robotic knee surgeries. While $M$-OR does not have real patients, it is more authentic as real clinicians perform the surgeries. We conduct four-fold cross-validation, partitioning by individual surgeries to ensure temporal separation between training and testing data. Following standard re-identification protocols [8], we adopt a probe-gallery evaluation framework where gallery sequences of known identities are compared against probe sequences of unknown identities.
@@ 3 Findings
GradCAM analysis reveals fundamentally different saliency heatmaps between datasets (see fig.˜1). On 4D-OR, CNNs consistently focus on the head and feet regions across all individuals, exploiting visible streetwear and distinctive footwear beneath surgical gowns. These artifacts provide strong identity signals within the dataset but represent spurious correlations that would fail in authentic clinical settings. In contrast, $M$-OR saliency maps reveal more diffuse, unfocused patterns lacking clear regions of attention. This dispersed activation suggests models struggle to identify reliable visual features when more distinguishing characteristics are not present due to standardized surgical attire.
@@
Quantitative evaluation confirms the limitations of RGB-based identification in realistic OR settings (see table˜1). While RGB and point cloud methods achieve comparable performance on 4D-OR (96.91% vs 95.95% macro accuracy, $p>0.05$), their effectiveness diverges dramatically on $M$-OR. Point cloud representations maintain high accuracy (85.75%) while RGB performance drops significantly to 73.23% ($p<0.05$). This 12% performance gap demonstrates that geometric features provide more reliable identity signals than appearance when in more realistic clinical settings without obvious simulation artefacts. The near-perfect RGB performance on 4D-OR further supports our hypothesis that models exploit dataset-specific visual artifacts rather than learning robust biometric features, suggesting the re-identification task is trivially solved. Cross-dataset evaluation amplifies these differences: models trained on one OR environment and tested on another show catastrophic RGB performance degradation (46.98% accuracy) while point cloud methods maintain reasonable transfer performance (65.92%). These results validate our hypothesis that shifting from appearance to geometric representations mitigates the impact of visual biases in the OR.

### v2

@@ 1 Introduction
Figure 1: RGB image excerpts and overlayed saliency maps generated with GradCAM [13] on the simulated datasets 4D-OR [12] and MM-OR [17]. 4D-OR’s limited realism and variety allow CNNs to identify individuals solely by their heads and shoes. In more realistic OR settings like $M$-OR, these features become less useful due to more homogenous attire. These discrepancies between different clinical environments can impede generalization.
@@ 2 Methods
We trained ResNet-50 models for person re-identification on two surgical datasets: the simulated 4D-OR [12] and the more realistic MM-OR [17]. Using GradCAM [13], we generated class activation maps to visualize which spatial regions the networks relied upon for identity discrimination. Models are trained on a supervised classification task of the individuals’ known identity.
@@
We evaluate our approach on two surgical datasets: the simulated 4D-OR [12] containing 5 individuals across 10 procedures, and the more authentic MM-OR [17] containing 13 individuals across 11 robotic knee surgeries. While MM-OR does not have real patients, it is more authentic as real clinicians perform the surgeries. We conduct four-fold cross-validation, partitioning by individual surgeries to ensure temporal separation between training and testing data. Following standard re-identification protocols [8], we adopt a probe-gallery evaluation framework where gallery sequences of known identities are compared against probe sequences of unknown identities.
@@ 3 Findings
GradCAM analysis reveals fundamentally different saliency heatmaps between datasets (see fig.˜1). On 4D-OR, CNNs consistently focus on the head and feet regions across all individuals, exploiting visible streetwear and distinctive footwear beneath surgical gowns. These artifacts provide strong identity signals within the dataset but represent spurious correlations that would fail in authentic clinical settings. In contrast, MM-OR saliency maps reveal more diffuse, unfocused patterns lacking clear regions of attention. This dispersed activation suggests models struggle to identify reliable visual features when more distinguishing characteristics are not present due to standardized surgical attire.
@@
Quantitative evaluation confirms the limitations of RGB-based identification in realistic OR settings (see table˜1). While RGB and point cloud methods achieve comparable performance on 4D-OR (96.91% vs 95.95% macro accuracy, $p>0.05$), their effectiveness diverges dramatically on MM-OR. Point cloud representations maintain high accuracy (85.75%) while RGB performance drops significantly to 73.23% ($p<0.05$). This 12% performance gap demonstrates that geometric features provide more reliable identity signals than appearance when in more realistic clinical settings without obvious simulation artefacts. The near-perfect RGB performance on 4D-OR further supports our hypothesis that models exploit dataset-specific visual artifacts rather than learning robust biometric features, suggesting the re-identification task is trivially solved. Cross-dataset evaluation amplifies these differences: models trained on one OR environment and tested on another show catastrophic RGB performance degradation (46.98% accuracy) while point cloud methods maintain reasonable transfer performance (65.92%). These results validate our hypothesis that shifting from appearance to geometric representations mitigates the impact of visual biases in the OR.